OpenStudy (anonymous):

Geniss HELP!!!!!!!!!!!!!!! What is the axis of symmetry for the function? y = x2 + 3x - 4 y = -2 y = -1.5 x = -2 x = -1. What are the values of a, b, and c in the function? List them in order with commas, such as 1,-2,3. y = -2x2 + 5x - 4

OpenStudy (anonymous):

i got the first anwser. its x=-1.5

OpenStudy (ash2326):

y=x^2+3x-4 y=x^2+3x+9/4-9/4-4 y=(x+3/2)^2-25/4 axis of symmetry is x=-3/2 you should check your options again 2nd question y=ax^2+bx+c\ y=-2x^2+5x-4 a=-2 b=5 c=-4

OpenStudy (ash2326):

you're correct